OpenStudy (anonymous):

Find the product. (2p^3)^3 · (3p^2)^2

OpenStudy (anonymous):

72*p^13

OpenStudy (anonymous):

binomial theorom

OpenStudy (anonymous):

multiply the top power

OpenStudy (anonymous):

i think something isnt right

OpenStudy (anonymous):

2p\[[(2p^2)^3] * [(3p^2)^2]\]

OpenStudy (anonymous):

2p\[2p^6*3p^4= 6p^(10)\]

OpenStudy (anonymous):

@Andresfon12 check the question well. it is (2p^3)^3 not (2p^2)^3

OpenStudy (anonymous):

adjaharrison u ar right right 9+4=13 2p^9∗3p^4=6p^(13)